Probability 3: Mastery Questions |
| Guidance Notes - To answer each question, click on the letter that corresponds to the correct answer. You then need to click the 'Check' button
to see if you are correct. If your answer is incorrect, you will be given a hint on how to get the correct answer. You then need to click the 'Try again' button to attempt
the question once more. |
| 1. |
Taro recorded the results of 30 spins of one of the spinners shown below.

Taro's results

Which is the most likely spinner Taro used?
Incorrect. Hint - Look at page 5 in Probability 3 unit.
|
| 2. |
The school hockey team had the following results last season: Won 13 games, drew 2 games and lost 5 games.
What is the relative frequency of losing?
Incorrect. Hint - Look at page 6 in Probability 3 unit.
|
| 3. |
The probability of winning a game is 0.6.
If you played the game 40 times, how many would you expect to win?
Incorrect. Hint - Look at page 6 in Probability 6 unit.
|
| 4. |
A box contains 50 cubes. If one cube is chosen at random the probability of getting a certain colour is given in the table below.
| Red |
Blue |
Green |
Yellow |
| 0.24 |
0.3 |
0.34 |
0.12 |
How many blue cubes are in the box?
Incorrect. Hint - what would 0.3 be written as fraction with the denominator 50.
|
The table below shows the probabilities of scoring 1, 2, 3 or 4 on a biased dice.
Questions 5 to 7 refer to this table. |
|
| X |
1 |
2 |
3 |
4 |
| P(X) |
0.4 |
0.15 |
0.25 |
0.2 |
|
| 5. |
What is P(X ≠ 3)?
Incorrect. Hint - P(X ≠ 3) = 1 − P(X = 3)
|
| 6. |
What is P(X < 3)?
Incorrect. Hint - P(x < 3) = P(X = 1) + P(X = 2)
|
| 7. |
What is P(1 ≤ X ≤ 3)?
Incorrect. Hint - P(1 ≤ X ≤ 3) = P(X = 1) + P(X = 2) + P(X = 3)
|
| 8. |
You roll two fair four sided dice and add up your scores.

What is the probability that you get a total score of 4?
Incorrect. Hint - Look at page 10 in Probability 3 unit.
|
| 9. |
If P(A) = 0.65, what is P(A')?
Incorrect. Hint - P(A') means P(X ≠ A).
|
| 10. |
Two fair pound coins are tossed together. What is the probability of getting two heads?
Incorrect. Hint - Look at page 10 in Probability 3 unit.
|
| 11. |

The Venn diagram above shows the number of students who played Netball and Rounders for the school team last week.
How many students did not play in the rounders team last week?
Incorrect. Hint - Look at page 20 in Probability 3 unit.
|
| For Questions 12 and 13, you will need to answer them using paper and pencil and then click the 'Check' button to see if you have done them correctly. |
| 12. |
Suitable sample space (1 mark)

(a) The probability of getting 3 heads = 18 (1 mark)
(b) The probability of getting 2 heads and a tail in any order = 38 (1 mark)
|
| 13. |
A running club has 50 members. 32 take part in road races (R) and 36 race on the track(T). 8 members do not take part in any running races.
This information is shown on the Venn diagram below.

a) The probability that a member runs in road races = 3250 or 1625 (1 mark)
b) The probability that a member does not run on the track = 1450 or 725 (1 mark)
c) The probability that a member runs on the road but not on the track = 650 or 325 (1 mark)
|
| ⮜Back to Year 7 Menu ⮜Back to Year 8 Menu |